You manage a pension fund that promises to pay out $10 million to its contributors in five years. You buy $7472582 worth of par-value bonds that make annual coupon payments of 6% and mature in five years. Right after you make the purchase, the interest rate on same-risk bonds decreases to 4.9%. If the rate does not change again and you reinvest the coupon payments that you receive in same-risk bonds, how much will you fall short of the money that you promised?
Answer #1

Money getting after 5 years=7472582*6%/4.9%*(1.049^5-1)+7472582=9945081.83964600

Falling short by 10*10^6-9945081.83964600=54918.16035400